OpenStudy (anonymous):

As an entrepreneur, there are going to be many decisions that you need to make, such as the price to charge your customers for your goods and services. You have just graduated from college and recently opened a specialty pizza restaurant. Based on surveys conducted in your area, you determine that it is feasible to sell your specialty pizzas for $15. The cost for making the pizzas includes a fixed cost of $55 and a labor cost of $4 per pizza. Establish an equation to determine revenue. Establish an equation to determine total cost. How many pizzas must be sold to break even.

OpenStudy (anonymous):

Full question is attached please help !!!!!!

OpenStudy (mertsj):

Revenue = number of pizzas times 15 R=15p C=55+4p Break even is when Revenue = Cost 15p=55+4p 11p=55 p=5 If you sell 5 pizzas, you will break even.

OpenStudy (mertsj):

Profit = Revenue - Cost The Revenue from 500 pizzas is R=15(500)=7500 The Cost to produce 500 pizzas is C= 55+4(500)=2055 Profit is 7500-2055=$5445 If you can make and sell 500 pizzas you will have a profit of $5445

OpenStudy (mertsj):

To make $1100 means your profit is 1100. Profit is Revenue - Cost Profit = 15p-(55+4p) Profit = 15p-55-4p Profit = 11p-55 1100=11p-55 1155=11p p=105 If you make and sell 105 pizzas, you will make $1100

OpenStudy (mertsj):

Profit = 1595 1595= 11p-55 1650=11p 150=p If you make and sell 150 pizzas you will make $1595. If you want to make more than $1595, you must make and sell more than 150 pizzas.